# Critical Metallicity of Cool Supergiant Formation. II. Physical Origin

### v1

@@ Abstract
This study investigates the physical origin of the critical metallicity ($Z_{\rm c}\sim 0.001$) required for the formation of cool supergiants, as revealed by stellar evolution models. Using model grids that vary in mass, metallicity, opacity, and nuclear reaction rates, we identify a threshold terminal-age main-sequence (TAMS) radius ($R_{\rm T}$) that determines whether a star of a given mass can evolve into the red supergiant (RSG) phase. Through stellar models and homology relations, we show that metallicity affects the TAMS radius via its influence on opacity and nuclear energy generation. By classifying the evolutionary pathways of supergiants, we demonstrate how TAMS radius, shaped by metallicity, decisively governs the post-main-sequence outcome: Stars with metallicities $Z<Z_{\rm c}$ exhibit TAMS radii smaller than $R_{\rm T}$ and proceed to advanced core helium or carbon burning while retaining compact envelopes, thereby preventing further expansion into the RSG regime. In contrast, stars with $Z>Z_{\rm c}$ have TAMS radii larger than $R_{\rm T}$ and expand into the stable RSG phase during core helium burning. The envelope radius at the onset of core helium burning is the key factor determining whether a star becomes a red or blue supergiant. Our results explain the origin of the critical metallicity and offer insight into the evolution of metal-poor massive stars in the early universe.
@@ 1 Introduction
In the first paper of this series (Ou et al., 2023, hereafter Paper I), we presented a grid of $\sim$2000 stellar models and identified a critical metallicity for the formation of cool supergiants (i.e., red or yellow), $Z_{\rm c}\sim 0.001\approx 0.1Z_{\odot}$ (For reference, $Z_{\odot}=0.0134$; Asplund et al., 2009). We found that massive stars with metallicities below this threshold ($Z<Z_{\rm c}$) typically do not evolve into cool supergiants during the core-helium (He) burning phase. This bifurcation in evolutionary paths has a significant impact on stellar mass loss, as cool supergiants exhibit considerably higher mass-loss rates than main-sequence stars. In this paper, we further explore the physical origin of this critical metallicity.
Our finding of a critical metallicity is consistent with various earlier stellar models, which likewise show that low-$Z$ stars tend to remain blue supergiants (BSGs) rather than evolving into red supergiants (RSGs) (Brunish & Truran, 1982; Arnett, 1991; Baraffe & El Eid, 1991; Brocato & Castellani, 1993; Hirschi, 2007; El Eid et al., 2009; Limongi, 2017). Similarly, in the Hertzsprung–Russell (HR) diagram of the recent Geneva stellar evolution models, stars with $Z=0.0004$ also remain blue during the core-He burning phase (Groh et al., 2019). These results suggest that the existence of a critical metallicity is a general outcome among various stellar evolution models.
@@
Although the metallicity dependence of RSG formation is consistent between models, the underlying physics behind the effect of metallicity remains uncertain. An early attempt to explain it was made by Ritossa (1996), who argued that the envelopes of low-metallicity stars cannot effectively trap radiation energy from central burning because of reduced opacity. Therefore, they cannot evolve into RSGs. This interpretation is based on a framework in which thermal imbalance within the envelope, governed by the core luminosity and the envelope opacity, drives the inflation of the stellar envelope towards red-giant or supergiant dimensions (Renzini, 1984; Renzini et al., 1992; Renzini & Ritossa, 1994; Renzini, 2023). However, this explanation for RSG formation based on heat absorption has been questioned by subsequent studies (Iben, 1993; Sugimoto & Fujimoto, 2000; Faulkner, 2005; Miller Bertolami, 2022).
@@
Understanding the metallicity effect on RSG formation is closely tied to the long-standing question: Why do stars become red supergiants? In our recent study (Ou & Chen, 2024), we revisited this question and identified the physical criteria that determine whether a star can evolve into the RSG phase. Building on those results, we now aim to investigate how metallicity influences RSG [^1] We briefly clarify our use of the terms ”red supergiant” and ”cool supergiant.” In Paper I, we adopted the term cool supergiant to describe massive stars that expand to $\gtrsim 1{,}000\,R_{\odot}$. We avoided the conventional term red supergiant because very massive stars (e.g., $\gtrsim 50\,M_{\odot}$) often maintain effective temperatures above $4{,}000\,{\rm K}$, placing them in the yellow rather than strictly defined red regime. Nevertheless, from a physical standpoint in terms of stellar radius, such stars do not differ intrinsically from classical red supergiants. Moreover, in this paper, our focus is mainly on stars with $\lesssim 30\,M_{\odot}$, which cool below $4{,}000\,{\rm K}$ and fulfill the criteria of red supergiants. Therefore, we adopt the more common term red supergiant throughout this work. formation to identify the physical origin of the critical metallicity using stellar models.
@@
In Section 2, we briefly describe the models employed in this study. In Section 3, we present a series of numerical experiments to investigate the metallicity effect on supergiant evolution. Section 4 identifies a threshold radius at the end of main sequence for RSG formation and explores its dependence on mass. To explore the connection between metallicity and this threshold radius, Section 5 studies how metallicity influences stellar radii. Section 6 classifies the evolutionary pathways for stellar models with different masses and metallicities, using the RSG formation criteria proposed by Ou & Chen (2024). In Section 7, we offer a physical interpretation of the threshold radius and critical metallicity based on the core properties. Finally, Section 8 discusses the broader implications of our findings and concludes the paper.
2 Models
@@
We begin with the grid of massive star models from Paper I, which spans initial masses from $11$ to $60\,M_{\odot}$ and metallicities from $Z=1\times 10^{-5}$ to $2\times 10^{-2}$. The models were computed using the Modules for Experiments in Stellar Astrophysics (MESA; Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023) version No. 10108. Each model evolves from the pre-main-sequence stage using the “basic” nuclear reaction network with eight isotopes, and then proceeds from the zero-age main sequence (ZAMS) using the “approx21 (Cr60)” network, which includes 21 isotopes. The models incorporate Henyey’s mixing-length theory, semi-convection, convective overshooting, Type 2 opacity, and mass loss. Details of the model setup are provided in Section 2 of Paper I.
@@
Following the analysis in Paper I, we adopt the $25\,M_{\odot}$ models as representative cases, as they exhibit typical evolutionary behaviors across different metallicities. Using these $25\,M_{\odot}$ models from the grid described above, we carry out a series of experiments on stellar models by modifying the opacity and nuclear reaction rates to examine their effects on supergiant evolution. These extended models are designed as numerical experiments to isolate the influence of specific physical parameters and are not intended to represent realistic stellar evolution.
Opacity, as a metallicity-dependent factor, strongly impacts the energy transport of the entire star. Adjusting opacity in stellar models requires careful consideration as opacity is coupled with the evolution of gas density, temperature, and metallicity. A robust method to modify opacity is by adjusting the value of $Z_{\rm base}$ in MESA, which specifies the base metallicity used to retrieve opacities from the opacity tables. Importantly, $Z_{\rm base}$ governs the opacity alone and does not represent the metallicity of the star. For example, a model with $Z=0.02$ and $Z_{\rm base}=0.001$ has a metallicity of 0.02 but evolves with an opacity comparable to that of a star with $Z=0.001$. This approach enables us to isolate the effects of opacity from other metallicity-dependent effects in stellar evolution. To avoid confusion in terminology, we denote the MESA input parameter $Z_{\rm base}$ as $\zeta_{\rm opacity}$ throughout the discussion in this paper.
@@
We also designed an experiment in which the CNO reaction rates are modified only during the shell H-burning phase in the post-main-sequence stage. Specifically, the adjustment is applied the adjustment is applied once the He core has formed and is limited to regions outside the core. By default in MESA, the He core is defined as the region where the H mass fraction satisfies $X_{\rm H}\leq 0.1$. For this experiment, we used the factor $\eta_{\rm shellCNO}$ to linearly scale the CNO reaction rates in the H burning shell.
In addition to the extensive grid of stellar models with varying masses and metallicities presented in Paper I, we have constructed four new distinct grids of MESA models spanning different parameter spaces. In total, these five grids contain 6,896 models. A summary of the model configurations is provided below and in Table 1.
@@
    Metallicity ($Z$) and opacity (set by $\zeta_{\rm opacity}$).
@@
    Hydrogen-burning rate (scaled by $\eta_{\rm CNO}$) and opacity (set by $\zeta_{\rm opacity}$).
@@
    Shell Hydrogen-burning rate (scaled by $\eta_{\rm shellCNO}$ for the burning shell) and opacity (set by $\zeta_{\rm opacity}$).
@@
    Helium-burning rate (scaled by $\eta_{3\alpha}$) and opacity (set by $\zeta_{\rm opacity}$).
Grid (a), as presented in Paper I, consists of realistic stellar models used to demonstrate the existence of a critical metallicity. On the other hand, the models in Grids (b)–(e) are not intended to represent realistic stars; instead, they act as experimental models to investigate the effects of specific stellar parameters. For comparison, we vary only two parameters in each grid of models, with all other parameters identical to those in Grid (a). Grids (b)--(e) are based on $25\,M_{\odot}$ stars, with Grids (c)–(e) further restricted to $Z=0.001$.
@@
Grid $M_{i}\ (M_{\odot})$ $Z$ $\zeta_{\rm opacity}$ $\eta_{\rm CNO}$ $\eta_{\rm shellCNO}$ $\eta_{3\alpha}$
(a) 11, 12, 13, $\ldots$, 59, 60 (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-5}$, (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-4}$, (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-3}$, (1.0, 1.1,, $(1.0,\,1.1,\,\ldots,\,2.0)\times 10^{-2}$ 1.0 1.0 1.0 1.0
(b) 25 (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-5}$, (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-4}$, (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-3}$, (1.0, 1.1,, $(1.0,\,1.1,\,\ldots,\,2.0)\times 10^{-2}$ (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-5}$, (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-4}$, (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-3}$, (1.0, 1.1,, $(1.0,\,1.1,\,\ldots,\,2.0)\times 10^{-2}$ 1.0 1.0 1.0
(c) 25 0.001 (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-5}$, (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-4}$, (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-3}$, (1, 2, 3, 4, $(1,\,2,\,3,\,4,\,5)\times 10^{-2}$ (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-3}$, (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-2}$, (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-1}$, (1, 2,, 10)1210(1,\,2,\,$\ldots$,\,10)( 1 , 2 , … , 10 ) 1.0 1.0
(d) 25 0.001 (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-5}$, (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-4}$, (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-3}$, (1, 2, 3, 4, $(1,\,2,\,3,\,4,\,5)\times 10^{-2}$ 1.0 (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-3}$, (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-2}$, (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-1}$, (1, 2,, 10)1210(1,\,2,\,$\ldots$,\,10)( 1 , 2 , … , 10 ) 1.0
(e) 25 0.001 (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-5}$, (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-4}$, (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-3}$, (1, 2, 3, 4, $(1,\,2,\,3,\,4,\,5)\times 10^{-2}$ 1.0 1.0 (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-3}$, (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-2}$, (1, 2,, $(1,\,2,\,\ldots,\,9)\times 10^{-1}$, (1, 2,, 10)1210(1,\,2,\,$\ldots$,\,10)( 1 , 2 , … , 10 )
Notes.
$M_{i}$: Initial stellar mass.
@@
$\zeta_{\rm opacity}$: Equivalent to the base metallicity ($Z_{\rm base}$) adopted for used for interpolating opacity in the tables, not to be confused with the star’s metallicity $Z$.
@@
3 Numerical Experiments on Opacity and Nuclear Reaction Rates
Metallicity affects stellar evolution through its impact on several key physical quantities, including opacity, nuclear burning rates, and mean molecular weight (e.g., El Eid et al., 2009). In this section, to disentangle the primary factors through which metallicity impacts evolution toward the supergiant phase, we conduct a series of experiments for $25\,M_{\odot}$ stars using Grid (b)–(e), as described in Section 2. In particular, we focus on the roles of opacity and nuclear reaction rates.
For each model in these grids, we assess whether it evolves into an RSG by calculating the maximum radius it reaches during the core He-burning phase, prior to the onset of core carbon (C) burning. This maximum radius is denoted as $R_{\rm max}$.
@@ 3 Numerical Experiments on Opacity and Nuclear Reaction Rates
Figure 1: Maximum stellar radius during the core He-burning stage ($R_{\rm max}$) for Grid (b), which includes models with an initial mass of $25\,M_{\odot}$, varying in metallicity ($Z$) and opacity parameter ($\zeta_{\rm opacity}$). The boundary between red and blue supergiants depends on $\zeta_{\rm opacity}$ in the metallicity range $Z\gtrsim 10^{-3}$, but becomes nearly insensitive to $\zeta_{\rm opacity}$ when $Z\lesssim 10^{-3}$.
@@
Figure 2: Maximum stellar radius during the core He-burning stage ($R_{\rm max}$) for Grid (c): models with an initial mass of $25\,M_{\odot}$ and metallicity $Z=0.001$, computed for varying values of the CNO-cycle reaction rate (scaled by $\eta_{\rm CNO}$) and opacity parameter ($\zeta_{\rm opacity}$). Both $\eta_{\rm CNO}$ and $\zeta_{\rm opacity}$ influence the evolutionary outcome, determining whether the star becomes a red or blue supergiant.
@@
Figure 3: Maximum stellar radius during the core He-burning stage ($R_{\rm max}$) for Grid (d): models with an initial mass of $25\,M_{\odot}$ and metallicity $Z=0.001$, computed for varying values of the CNO-cycle reaction rate during the shell H-burning stage (scaled by $\eta_{\rm shellCNO}$) and opacity parameter ($\zeta_{\rm opacity}$). For comparison, the parameter $\eta_{\rm CNO}$ used in Figure 2 is applied throughout all evolutionary stages, whereas $\eta_{\rm shellCNO}$ in this figure is applied only during the shell H-burning stage. While $R_{\rm max}$ exhibits some scatter, the overall trend indicates that $\eta_{\rm shellCNO}$ has little impact on the supergiant outcome.
@@
Figure 4: Maximum stellar radius during the core He-burning stage ($R_{\rm max}$) for Grid (e): models with an initial mass of $25\,M_{\odot}$ and metallicity $Z=0.001$, computed for varying values of the triple-alpha reaction rate (scaled by $\eta_{3\alpha}$) and opacity parameter ($\zeta_{\rm opacity}$). Increasing $\eta_{3\alpha}$ does not promote evolution toward the RSG phase and may even suppress it in the low-opacity regime ($\zeta_{\rm opacity}\lesssim 10^{-3}$).
3.1 Grid (b): Metallicity and Opacity
We begin by examining the effect of opacity on envelope expansion by constructing a grid of stellar models, Grid (b), with varying metallicities and opacities.
Figure 1 shows $R_{\rm max}$ for each model in Grid (b). The models that reach RSG dimensions ($\log(R/R_{\odot})\gtrsim 3$) are clearly separated from those that remain within BSG dimensions ($\log(R/R_{\odot})\lesssim 2$). For $Z\lesssim 10^{-3}$, the boundary dividing these two areas is vertical, indicating that opacity, controlled by $\zeta_{\rm opacity}$, has little influence on the supergiant outcome in this low-$Z$ range. Only when $Z\gtrsim 10^{-3}$ does opacity begin to significantly affect whether the star becomes an RSG or remains a BSG. Notably, even with extremely low $\zeta_{\rm opacity}$ (a few times $10^{-5}$), stars with $Z\gtrsim 0.003$ can still expand to RSG dimensions. This suggests that additional metallicity-dependent factors, beyond opacity alone, play a role in determining the supergiant’s fate.
3.2 Grid (c): Hydrogen Burning Rates and Opacity
In addition to opacity, metallicity also significantly influences nuclear burning within a star. In our second experiment, Grid (c), we examine the combined effects of these two factors by varying both the opacity parameter $\zeta_{\rm opacity}$ and the linear scaling factor for the reaction rates of the CNO cycle, $\eta_{\rm CNO}$. In this experiment, $\eta_{\rm CNO}$ is applied consistently throughout the entire stellar lifetime, scaling the reaction rates of both core H burning during the main-sequence phase and shell H burning during the post-main-sequence evolution.
Figure 2 shows $R_{\rm max}$ for each model in Grid (c). A clear boundary separates models that evolve into RSGs from those that remain as BSGs. In general, both higher opacity and higher CNO reaction rates favor the transition to RSGs. When the opacity is low, a larger CNO rate is required for the star to expand to RSG radii. In particular, even models with low opacity with $\zeta_{\rm opacity}$ on the order of a few times $10^{-5}$ can still evolve into RSGs if CNO burning is sufficiently efficient (i.e., $\eta_{\rm CNO}\geq 3$).
3.3 Grid (d): Shell Hydrogen Burning Rate and Opacity

### v2

@@ Abstract
This study investigates the physical origin of the critical metallicity required for the formation of cool supergiants, as revealed by stellar evolution models. Using grids of stellar models, we show that the terminal-age main-sequence (TAMS) radius, $R_{\rm TAMS}$, defines a threshold that determines whether a star of a given mass can evolve into the red supergiant (RSG) phase. Metallicity influences the supergiant outcome because it modifies $R_{\rm TAMS}$ through its effects on opacity and nuclear energy generation, as demonstrated by our stellar models and dimensional analysis based on homology relations. The value of $R_{\rm TAMS}$ sets the initial radius for post-main-sequence expansion and therefore controls the envelope radius reached at subsequent core-evolution stages. Higher-metallicity stars develop larger $R_{\rm TAMS}$ and rapidly expand into the stable RSG regime during core helium burning. In contrast, lower-metallicity stars have smaller $R_{\rm TAMS}$ and advance to more evolved core helium or carbon-burning stages while retaining compact envelopes, thereby preventing expansion into the RSG regime during core helium burning. Our results explain the origin of the critical metallicity and offer insight into the evolution of metal-poor massive stars in the early universe.
@@ I Introduction
In the first paper of this series (Ou et al., 2023, hereafter Paper I), we presented a grid of $\sim$2000 non-rotating massive star models and identified a critical metallicity at around $Z\sim 0.001$ for the formation of cool supergiants (i.e., red or yellow). We found that stars with metallicities below this threshold typically do not evolve into cool supergiants during the core-helium (He) burning phase. Our finding is consistent with various earlier stellar models, which likewise show that low-$Z$ stars tend to remain blue supergiants (BSGs) rather than evolving into red supergiants (RSGs) (Brunish & Truran, 1982; Arnett, 1991; Baraffe & El Eid, 1991; Brocato & Castellani, 1993; Hirschi, 2007; El Eid et al., 2009; Limongi, 2017; Groh et al., 2019), but we further identified a well-defined metallicity threshold for this transition. This bifurcation in evolutionary paths has a significant impact on stellar mass loss, as RSGs exhibit considerably higher mass-loss rates than main-sequence stars and BSGs. In this paper, we investigate the physical origin of this critical metallicity.
@@
An early attempt to explain the metallicity dependence of RSG formation was made by Ritossa (1996), who argued that the envelopes of low-metallicity stars cannot effectively trap radiation energy from central burning because of reduced opacity, and therefore cannot evolve into RSGs. This interpretation is based on a framework in which thermal imbalance within the envelope, governed by the core luminosity and the envelope opacity, drives the inflation of the stellar envelope to red-giant or supergiant dimensions (Renzini, 1984; Renzini et al., 1992; Renzini & Ritossa, 1994; Renzini, 2023). However, this explanation for RSG formation based on heat absorption has been shown to be inconsistent with subsequent studies (Iben, 1993; Sugimoto & Fujimoto, 2000; Faulkner, 2005; Miller Bertolami, 2022; Ou & Chen, 2024).
@@
Understanding the metallicity effect on RSG formation is closely tied to the long-standing question: Why do stars become RSGs? Our recent study (Ou & Chen, 2024) proposed that envelope expansion toward the RSG phase follows the refined "mirror principle," under which there exist two pathways for stars to evolve into RSGs – direct expansion during core contraction, and continued expansion after core contraction governed by nuclear energy generation rates. Building on those results, we now aim to investigate how metallicity influences RSG [^1] We briefly clarify our use of the terms ”red supergiant” and ”cool supergiant.” In Paper I, we adopted the term cool supergiant to describe massive stars that expand to $\gtrsim 1{,}000\,R_{\odot}$. We avoided the conventional term red supergiant because very massive stars (e.g., $\gtrsim 50\,M_{\odot}$) often maintain effective temperatures above $4{,}000\,{\rm K}$, placing them in the yellow rather than strictly defined red regime. Nevertheless, from a physical standpoint in terms of stellar radius, such stars do not differ intrinsically from classical red supergiants. Moreover, in this paper, our focus is mainly on stars with $\lesssim 30\,M_{\odot}$, which cool below $4{,}000\,{\rm K}$ and fulfill the criteria of red supergiants. Therefore, we adopt the more common term red supergiant throughout this work. formation to identify the physical origin of the critical metallicity using stellar models.
@@
In Section II, we briefly describe the models employed in this study. In Section III, we present a series of numerical experiments to investigate the metallicity effect on supergiant evolution. Section IV identifies a threshold radius at the end of main sequence for RSG formation and explores its dependence on mass. To explore the connection between metallicity and this threshold radius, Section V demonstrates how metallicity influences stellar radii. In Section VI, we offer a physical interpretation of the threshold radius and critical metallicity based on He core evolution. Finally, Section VII discusses the broader implications and concludes the paper.
II Models
@@
We begin with the grid of massive star models from Paper I, hereafter Grid (a), which spans initial masses from $11$ to $60\,M_{\odot}$ and metallicities from $Z=1\times 10^{-5}$ to $2\times 10^{-2}$. The models were computed using the Modules for Experiments in Stellar Astrophysics (MESA; Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023) version No. 10108. Each model evolves from the pre-main-sequence stage using the "create_zams" module with the “basic” nuclear reaction network, and then proceeds from the zero-age main sequence (ZAMS) using the “approx21 (Cr60)” network, which includes 21 isotopes. The models adopt Henyey’s mixing-length theory with a mixing-length parameter of $\alpha_{\rm MLT}=1.5$. Semi-convection is included with an efficiency of $\alpha_{\rm sc}=0.01$. Exponential overshooting is implemented following Herwig (2000), with $f_{\rm ov}=0.001$ in non-burning and H-burning regions, $f_{\rm ov}=0$ in He- and metal-burning regions, and $f_{0,\,\rm ov}=0.0005$ in all regions. Further details of the model setup are provided in Section 2 of Paper I.
@@
Using the $25\,M_{\odot}$ models from Grid (a), we carry out a series of controlled experiments in which we individually modify several metallicity-dependent factors, such as opacity and nuclear reaction rates, to assess their impact on supergiant evolution. These numerical experiments are designed to separate the influence of specific physical parameters and are not intended to represent realistic stellar evolution.
To control opacity systematically, we vary the MESA parameter $Z_{\rm base}$, which specifies the base metallicity used to query the opacity tables. This parameter affects the opacity alone and is independent of the actual stellar metallicity. For example, a model with $Z=0.02$ and $Z_{\rm base}=0.001$ retains its true metallicity of 0.02 but evolves with opacities comparable to that of a star with $Z=0.001$. This method helps to separate the effects of opacity from other metallicity-dependent influences, such as changes in nuclear energy generation and the mean molecular weight. To avoid confusion, we refer to $Z_{\rm base}$ as $\zeta_{\kappa}$ throughout this paper.
@@
We also perform an experiment in which the CNO reaction rates are modified only during the post-main-sequence shell H-burning phase. The adjustment is applied after the He core forms and is restricted to regions outside the core, which MESA defines as zones with $X_{\rm H}>0.1$. In this experiment, we linearly scale the CNO reaction rates in the H burning shell by a factor $\eta_{\rm shellCNO}$.
In total, we construct five grids of MESA models covering different parameter spaces, comprising 6,896 models:
@@
    Metallicity ($Z$) and opacity (set by $\zeta_{\kappa}$).
@@
    Hydrogen-burning rate (scaled by $\eta_{\rm CNO}$) and opacity (set by $\zeta_{\kappa}$).
@@
    Shell Hydrogen-burning rate (scaled by $\eta_{\rm shellCNO}$ for the burning shell) and opacity (set by $\zeta_{\kappa}$).
@@
    Helium-burning rate (scaled by $\eta_{3\alpha}$) and opacity (set by $\zeta_{\kappa}$).
Grids (b)–(e) serve as controlled experiments for separating the effects of specific stellar parameters. In each grid, only two parameters are varied, while all others are kept identical to Grid (a). Grids (b)–(e) are based on $25\,M_{\odot}$ models, with Grids (c)–(e) further restricted to $Z=0.001$. The full set of model parameters is summarized in Table 1.
@@
Grid $M_{i}\ (M_{\odot})$ $Z$ $\zeta_{\kappa}$ $\eta_{\rm CNO}$ $\eta_{\rm shellCNO}$ $\eta_{3\alpha}$
[…]
Note. — $M_{i}$: Initial stellar mass.
@@
$\zeta_{\kappa}$: Equivalent to the base metallicity ($Z_{\rm base}$) adopted for interpolating opacity in the tables, and should not be confused with the stellar metallicity $Z$.
@@
III Numerical Experiments on Opacity and Nuclear Reaction Rates
Metallicity influences stellar evolution through its effects on opacity, nuclear burning rates, and the mean molecular weight (e.g., El Eid et al., 2009). In this section, we use Grids (b)–(e) to separate the roles of opacity and nuclear reaction rates in determining whether a star evolves toward the RSG phase. For each model in these grids, we evaluate its evolutionary outcome by computing the maximum radius attained during core He burning—before the onset of core carbon (C) burning—which we denote as $R_{\rm max}$. The values of $R_{\rm max}$ are plotted in grid plots, from which we observe the division between models that reach RSG dimensions ($\log(R/R_{\odot})\gtrsim 3$) and those that remain in the BSG regime ($\log(R/R_{\odot})\lesssim 2$). We examine whether the boundary between these two regimes aligns along any threshold of opacity or nuclear reaction rate parameters.
@@ III Numerical Experiments on Opacity and Nuclear Reaction Rates
Figure 1: Maximum stellar radius during the core He-burning stage ($R_{\rm max}$) for Grid (b), which includes models with an initial mass of $25\,M_{\odot}$, varying in metallicity ($Z$) and opacity parameter ($\zeta_{\kappa}$). The boundary between red and blue supergiants depends on $\zeta_{\kappa}$ in the range $\zeta_{\kappa}\gtrsim 10^{-3}$, but becomes nearly insensitive to $\zeta_{\kappa}$ when $\zeta_{\kappa}\lesssim 10^{-3}$.
@@
Figure 2: Maximum stellar radius during the core He-burning stage ($R_{\rm max}$) for Grid (c): models with an initial mass of $25\,M_{\odot}$ and metallicity $Z=0.001$, computed for varying values of the CNO-cycle reaction rate (scaled by $\eta_{\rm CNO}$) and opacity parameter ($\zeta_{\kappa}$). Both $\eta_{\rm CNO}$ and $\zeta_{\kappa}$ influence the evolutionary outcome, determining whether the star becomes a red or blue supergiant.
@@
Figure 3: Maximum stellar radius during the core He-burning stage ($R_{\rm max}$) for Grid (d): models with an initial mass of $25\,M_{\odot}$ and metallicity $Z=0.001$, computed for varying values of the CNO-cycle reaction rate during the shell H-burning stage (scaled by $\eta_{\rm shellCNO}$) and opacity parameter ($\zeta_{\kappa}$). For comparison, the parameter $\eta_{\rm CNO}$ used in Figure 2 is applied throughout all evolutionary stages, whereas $\eta_{\rm shellCNO}$ in this figure is applied only during the shell H-burning stage. While $R_{\rm max}$ exhibits some scatter, the overall trend indicates that $\eta_{\rm shellCNO}$ has little impact on the supergiant outcome.
@@
Figure 4: Maximum stellar radius during the core He-burning stage ($R_{\rm max}$) for Grid (e): models with an initial mass of $25\,M_{\odot}$ and metallicity $Z=0.001$, computed for varying values of the triple-alpha reaction rate (scaled by $\eta_{3\alpha}$) and opacity parameter ($\zeta_{\kappa}$). Increasing $\eta_{3\alpha}$ does not promote evolution toward the RSG phase and may even suppress it in the low-opacity regime ($\zeta_{\kappa}\lesssim 10^{-3}$).
III.1 Grid (b): Metallicity and Opacity
Grid (b) varies the metallicity and opacity for $25\,M_{\odot}$ models to test how opacity affects the supergiant outcome. The resulting values of $R_{\rm max}$ are shown in Figure 1, which reveals a clear separation between RSG and BSG outcomes (red and blue, respectively). When $\zeta_{\kappa}\gtrsim 10^{-3}$, opacity can shift the metallicity threshold for RSG formation, with higher opacities allowing RSGs to form at lower $Z$. However, for $\zeta_{\kappa}\lesssim 10^{-3}$, the boundary between RSG and BSG models becomes vertical, indicating that opacity has little influence on the evolutionary outcome in this regime. Notably, even with extremely low $\zeta_{\kappa}$ (a few times $10^{-5}$), stars with $Z\gtrsim 0.003$ still expand to RSG dimensions. This implies that opacity may not be the only metallicity-dependent factor that determines RSG formation.
III.2 Grid (c): Hydrogen Burning Rate and Opacity
Metallicity also strongly affects nuclear burning. Using Grid (c), we examine the combined influence of two metallicity-dependent factors—opacity (set by $\zeta_{\kappa}$) and the hydrogen-burning rate (scaled by $\eta_{\rm CNO}$ and applied to both core H burning on the main sequence and shell H burning thereafter). As shown in Figure 2, the boundary between the red (RSG) and blue (BSG) regions indicates that both higher opacity and higher H burning rate promote the transition to RSGs. Even when the opacity is low (e.g., $\zeta_{\kappa}$ of a few $\times 10^{-5}$), the star can still evolve into an RSG if H burning is sufficiently efficient (i.e., $\eta_{\rm CNO}\geq 3$).
III.3 Grid (d): Shell Hydrogen Burning Rate and Opacity